# DiffOPF: Diffusion Solver for Optimal Power Flow

### v1

@@ I Introduction
The seminal works in [6, 7, 8] have laid the foundations of deep learning solutions to OPF. Given the load input, the solver in [6] yields an efficient warm start for a non-convex load-flow problem, significantly reduces the computational burden. Since then, substantial research efforts have focused on improving the performance of load-to-dispatch mappings, with emphasis on operational and security constraints satisfaction [7, 9, 10, 11], integrating extra domain knowledge [12], resolving uncertainties [13], compressing the mapping [14, 15], and developing topology-informed architectures for deep learning OPF solvers [16, 17]. The interested reader is referred to [18] for a comprehensive list of contributions in this area.
@@
However, the inherent complexities of the load-to-dispatch OPF mapping pose practical challenges for single-valued deep learning OPF solvers, as depicted in Fig. 1. First, the underlying OPF mapping admits multiple global and local solutions [19], which cannot be fully captured by single-valued solvers unless provided by extra information. This problem has been addressed by work in [20], where the authors augmented the feature space with initialization for power injection and voltage variables. Second, power system operations are influenced by controlled and uncontrolled disturbances, both contributing to the stochastic nature of each operational record. The former disturbances relate to constantly changing transformer tap settings, shunt and topology switching, and corrective actions. The latter disturbances arise from weather- and market-induced variability of OPF parameters, such as temperature-sensitive admittances, cost functions affected by fuel prices, and market strategies among others. While both disturbances can be incorporated in the design of deep learning OPF solvers (e.g., by inflating the feature space [20] or integrating the grid topology [16, 17]), this task is infeasible to account for all types of disturbances at once. Moreover, associating each operational record with specific realizations of disturbances may not be possible due to lack of data.
In this paper, we develop a data-driven solver that addresses the multi-valued nature of the load-to-dispatch mapping in OPF. The mapping is abstracted by a diffusion model, termed DiffOPF, whose output is the conditional distribution of generator setpoints given load inputs. Sampling from the conditional distribution allows system operators to explore various warm starts for the downstream load-flow problem, and thereby assess the associated dispatch cost and constraint satisfaction outcomes. Our contributions include:
@@
    The conditional sampling enables system operators to quantify uncertainty in dispatch setpoints and generate multiple statistically credible warm starts for a given load level. Each sample provides distinct cost and constraint satisfaction outcomes when used in load-flow problems, allowing operators to select solutions that best align with current operation and risk tolerance. Thus, DiffOPF improves on single-valued deep learning OPF solvers in warm start applications.
@@
More broadly, our work aligns with the emerging concept of decision flow for sampling decisions from target distributions guided by prior samplers [24]. Preliminary applications of decision flow with custom diffusion processes to unit commitment appear in [25]. Here, however, we leverage established DDPM [21] and score-based diffusion models [22] rather than developing custom diffusion processes. Prior work [26] demonstrated that DDPM suits OPF applications, as it enables control of the output distribution to satisfy power flow equations and grid constraints for the synthesis of power flow datasets. Although this extension is natural for data-driven OPF solvers [7, 9, 11], this paper establishes the foundational principles of DiffOPF for warm start applications, relegating full constraint satisfaction to future work.
@@ II Problem Setup
Sampling from the conditional distribution produces statistically credible warm start solutions for the downstream load flow problem. An overview of this setup is given in Fig. 2.
@@ III-B Diffusion Models
The training procedure is summarized in Alg. 1 [21]. Once the neural network is trained, new OPF data points can be generated by first estimating the clean sample $\hat{\mathbf{x}}_{0}$ at each step $t$ using the noise predictor $\bm{s}_{\theta}(\mathbf{x}_{t},t)$:
@@ IV-A DiffOPF Solver
where $\lambda=1/\sigma_{g}^{2}$ is the scale parameter for guidance, which can be empirically tuned for performance. Algorithm 3 incorporates this guidance term into sampling, enforcing consistency of dispatch setpoints with observed demands.
Algorithm 3 : Sampling OPF warm starts with guidance
[…]
2:for t=T1t=T-1 to 0 do
5:  $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$
7:return $(\mathbf{p}_{g},\mathbf{q}_{g})\subset\mathbf{x}_{0}$
@@ IV-B Sample Complexity of DiffOPF
In contrast, the proposed DiffOPF in Alg. 3 aims to approximate the full conditional distribution $p(\mathbf{p}_{g},\mathbf{q}_{g}|\mathbf{p}_{d},\mathbf{q}_{d})$. Formally, given demand $(\mathbf{p}_{d},\mathbf{q}_{d})$, the solver generates candidate solutions by sampling the conditional distribution $p(\mathbf{p}_{g},\mathbf{q}_{g}\mid\mathbf{p}_{d},\mathbf{q}_{d})$. That is, rather than reducing the solution space of the OPF to the mean values, DiffOPF provides a distributional characterization of dispatch setpoints for given demand.
@@
as the probability that a single sample is $\varepsilon$-close to the locally OPF solution $c^{\star}$. By independence, the probability that none of the $M$ samples is $\varepsilon$-close to $c^{\star}$ is $(1-p_{\varepsilon})^{M}$. Taking the complement gives the probability that at least one sample has this property:
@@
V Load Flow Problem and Evaluation of DiffOPF Warm-Starts
We propose DiffOPF for warm-start applications, where the solver provides an initialization for a downstream load-flow problem. In this regard, for each generated sample $(\mathbf{p}_{g},\mathbf{q}_{g})$ we solve the following load-flow problem [7]:
@@ V Load Flow Problem and Evaluation of DiffOPF Warm-Starts
    defined as the cost difference between the load-flow solution and the AC-OPF from the historical dataset.
@@ VI-A Experimental Setup
Working with DiffOPF, for each load instance, we sample 250 dispatch setpoints as warm starts for the load-flow problem (20), and then report the results for the best sample in terms of each metric in (21). The sampling time across all test cases does not exceed $2.3$ seconds per DiffOPF sample. When implemented on distributed computing nodes, all samples can be obtained at once, taking $2.3$ seconds in total.
@@ VI-C Tests Across Benchmark Power Systems
Table I summarizes the outcomes of the load-flow problem (20) that receives warm starts from DeepOPF and DiffOPF models for selected benchmark power systems. The table provides statistics across 100 load instances for the three metrics defined in (21), including the mean values over 100 load instances and the mean of 10 worst-case instances, represented by the conditional value at risk at the 10th quantile $\text{CVaR}_{10\%}$.
@@ VI-D Sample Complexity
We validate the sample complexity requirement of DiffOPF introduced in Sec. IV-B, and provide the number of samples required to obtain at least one $\varepsilon$-close solution at different confidence levels $\delta$ for the smallest PJM $5$-bus and the largest IEEE $118$-bus test cases. To compute $p_{\varepsilon}$, we consider the same $100$ test load instances as in Sec. VI-C. For each load instance, we draw $250$ samples from DiffOPF and count how many of them are $\varepsilon$-close to the solution provided by conventional interior-point method. We then take the average over all load instances and divide this average by the total number of samples to obtain the empirical probability $p_{\varepsilon}$.
Table II shows the theoretical lower bound on the required number $M$ of samples to have at least one $\varepsilon$-close solution, and compares it with the actual number $N_{\varepsilon}$ of generated samples within the $\varepsilon$ distance. Note, we present the distance $\varepsilon$ in percentage, as the cost structure in the two systems is different.
The lower bound appears conservative for the small test case PJM $5$-bus, where we have relatively large $N_{\varepsilon}$ with respect to $M$. However, for the large test case IEEE $118$-bus, the lower bound is tighter, and $N_{\varepsilon}$ is relatively small. As $\varepsilon$ increases, fewer samples are needed to obtain an $\varepsilon$-close solution. Even for small distance like $\varepsilon=0.5\%$, DiffOPF only requires a modest number of samples, not exceeding $269$.
@@ VII Conclusion
For future work, we intend to address the runtime gap between DiffOPF, which takes at most $2.3$ seconds for IEEE $118$-bus system, and single-valued deep learning OPF solvers with millisecond runtimes. We will focus on accelerating DiffOPF via diffusion distillation techniques, which compress sampling into a few denoising steps [30]. To extend beyond warm start applications, we will also build on [26] to incorporate constraints within the diffusion process.

### v2

@@ I Introduction
The seminal works in [34, 7, 20] have laid the foundations of deep learning solutions to OPF. Given the load input, the solver in [34] yields an efficient warm start for a non-convex OPF problem, significantly reduces the computational burden. Since then, substantial research efforts have focused on improving the performance of load-to-dispatch mappings, with emphasis on operational and security constraints satisfaction [7, 21, 30, 17], integrating extra domain knowledge [25], resolving uncertainties [8], compressing the mapping [24, 13], and developing topology-informed architectures for deep learning OPF solvers [15, 31]. The interested reader is referred to [18] for a comprehensive list of contributions in this area.
@@
However, the inherent complexities of the load-to-dispatch OPF mapping pose practical challenges for single-valued deep learning OPF solvers, as depicted in Fig. 1. First, the underlying OPF mapping admits multiple global and local solutions [19], which cannot be fully captured by single-valued solvers unless provided by extra information. This problem has been addressed by work in [22], where the authors augmented the feature space with initialization for power injection and voltage variables. Second, power system operation is influenced by controlled and uncontrolled disturbances, both contributing to the stochastic nature of each operational record. The former disturbances relate to constantly changing transformer tap settings, shunt and topology switching, and corrective actions. The latter disturbances arise from weather- and market-induced variability of OPF parameters, such as temperature-sensitive admittances, cost functions affected by fuel prices, and market strategies among others. While some sources of variability, such as network parameters or topology, can in principle be incorporated in the design of deep learning OPF solvers (e.g., by expanding the feature space [22] or integrating grid topology [15, 31]), accounting for all operational factors in this way is generally infeasible. Moreover, associating each operational record with the precise realization of these factors is often not possible due to incomplete or unavailable data.
In this paper, we develop a data-driven solver that addresses the multi-valued nature of the load-to-dispatch mapping in OPF. The mapping is abstracted by a diffusion model, termed DiffOPF, whose output is the conditional distribution of generator setpoints given load inputs. Sampling from the conditional distribution allows system operators to generate warm-start points for downstream power flow problems, enabling efficient exploration of dispatch options and cost vs. constraint satisfaction trade-offs. Our contributions include:
@@
    The conditional sampling enables system operators to quantify uncertainty in dispatch setpoints and generate multiple statistically credible warm starts for a given load level. Each sample provides distinct cost and constraint satisfaction outcomes, allowing operators to select solutions that best align with current operation and risk tolerance. Thus, DiffOPF improves on single-valued deep learning OPF solvers in warm start applications.
@@
More broadly, our work aligns with the emerging concept of decision flows, which aims to sample decisions from target distributions guided by prior samplers [3]. Preliminary applications of decision flows with custom diffusion processes to unit commitment appear in [6]. In this work, however, we leverage established DDPM [9] and score-based diffusion models [26] rather than developing custom diffusion processes.
While other generative models such as GANs and VAEs have been explored in power systems, their applications have primarily focused on generating synthetic data, such as load or renewable generation profiles [5, 23], rather than addressing inverse problems arising in system operation. Moreover, in GAN-based approaches, constraints are typically incorporated through penalization in the training objective, which can exacerbate training instability and mode collapse [32, 11]. In contrast, diffusion models provide a principled mechanism for incorporating constraints during sampling. This capability forms the foundation of DiffOPF, enabling the generation of dispatch solutions conditioned on demand levels, which act as structural constraints in the inverse problem.
Prior work [10] demonstrated that DDPMs are well suited for OPF-related applications, enabling control of generated samples to satisfy power flow constraints in the synthesis of power flow datasets. Although extending such constraint enforcement to data-driven OPF solvers is a natural next step [7, 21, 17], this paper focuses on establishing the foundational principles of DiffOPF for warm-start applications, leaving full constraint satisfaction for future work.
@@ II Problem Setup
Sampling from the conditional distribution produces statistically credible warm start solutions for the downstream power flow problem. An overview of this setup is given in Fig. 2.
@@ III-B Diffusion Models
The training procedure is summarized in Alg. 1 [9]. Once the neural network is trained, new OPF data points can be generated by first estimating the clean sample $\hat{\mathbf{x}}_{0}$ at each step $t$ using the score predictor $\bm{s}_{\theta}(\mathbf{x}_{t},t)$:
@@ IV-A DiffOPF Solver
Algorithm 3 : Sampling OPF warm starts with guidance
[…]
2:for t=T1t=T-1 to 0 do
5:  $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$
7:return $(\mathbf{p}_{g},\mathbf{q}_{g})\subset\mathbf{x}_{0}$
@@
where $\lambda=1/\sigma_{g}^{2}$ is the scale parameter for guidance, which can be empirically tuned for performance.
Algorithm 3 incorporates this guidance term into the reverse diffusion process, steering the sampling trajectory toward dispatch solutions consistent with the observed demands. However, the sampling procedure remains stochastic. In particular, two sources of randomness are present: the initial noise $\mathbf{x}_{T}$ used to start the reverse process at $t=T$, and the noise $\mathbf{z}$ injected at each reverse step. These stochastic components lead to different sampling trajectories, allowing the model to explore multiple high-probability regions of the learned conditional distribution. Hence, when historical data contain multiple dispatch patterns associated with similar load levels, different noise realizations can produce diverse candidate dispatch solutions conditioned on the same load.
In the present work, the guidance term only enforces consistency with the specified demand levels and does not explicitly incorporate the OPF constraints during sampling. Therefore, the generated samples are not guaranteed to be fully feasible OPF solutions. Instead, they should be interpreted as candidate dispatch points that can serve as warm-starts for conventional OPF solvers.
Beyond warm-start applications, learning the conditional distribution of generator dispatch given load levels enables several additional use cases. For example, it can help characterize the generation patterns associated with a given load realization, generate diverse candidate operating points for contingency and security assessment, and support uncertainty-aware operational studies. These capabilities arise from modeling a distribution of dispatch solutions rather than producing a single deterministic prediction.
@@ IV-B Sample Complexity of DiffOPF
In contrast, the proposed DiffOPF in Alg. 3 aims to approximate the full conditional distribution $p(\mathbf{p}_{g},\mathbf{q}_{g}|\mathbf{p}_{d},\mathbf{q}_{d})$. Formally, given demand $(\mathbf{p}_{d},\mathbf{q}_{d})$, the solver generates candidate solutions by sampling the conditional distribution $p(\mathbf{p}_{g},\mathbf{q}_{g}\mid\mathbf{p}_{d},\mathbf{q}_{d})$. If historical OPF records contain multiple dispatch solutions associated with similar load levels, these appear as distinct modes in the conditional distribution $p(\mathbf{p}_{g},\mathbf{q}_{g}|\mathbf{p}_{d},\mathbf{q}_{d})$. Since diffusion models learn the full data distribution, conditional sampling allows the model to generate samples around these modes, thus providing a distributional characterization of dispatch setpoints rather than reducing the solution space to mean values as in single-valued deep learning OPF solvers.
@@
as the probability that a single sample is $\varepsilon$-close to the locally OPF solution $c^{\star}$. By independence, the probability that none of the $M$ samples is $\varepsilon$-close to $c^{\star}$ is $(1-p_{\varepsilon})^{M}$. The independence assumption is justified because each sample is generated by an independent run of the diffusion sampler. Due to Alg. 3, the reverse diffusion process is initialized from an independent Gaussian noise realization and evolves with independent stochastic updates.
Taking the complement gives the probability that at least one sample has this property:
@@
V Evaluation of DiffOPF Warm-Starts
We propose DiffOPF for warm-start applications, where the solver provides an initialization for a downstream projection problem. In this regard, for each generated sample $(\mathbf{p}_{g},\mathbf{q}_{g})$ we solve the following problem [7]:
@@ V Evaluation of DiffOPF Warm-Starts
    defined as the cost difference between the projection (20) solution and the dispatch from the historical dataset.
@@ VI-A Experimental Setup
Working with DiffOPF, for each load instance, we sample 250 dispatch setpoints for the projection problem (20), and then report the results for the best sample in terms of each metric in (21). The sampling time across all test cases does not exceed $2.3$ seconds per DiffOPF sample. When implemented on distributed computing nodes, all samples can be obtained at once, taking $2.3$ seconds in total.
@@ VI-C Tests Across Benchmark Power Systems
Table I summarizes the outcomes of the projection problem (20) that receives warm starts from DeepOPF and DiffOPF models for selected benchmark power systems. The table provides statistics across 100 load instances for the three metrics defined in (21), including the mean values over 100 load instances and the mean of 10 worst-case instances, represented by the conditional value at risk at the 10th quantile $\text{CVaR}_{10\%}$.
@@ VI-D Sample Complexity
We validate the sample complexity requirement of DiffOPF introduced in Sec. IV-B, and report the number of samples required to obtain at least one $\varepsilon$-close solution at different confidence levels $\delta$ for the smallest PJM $5$-bus and the largest IEEE $118$-bus test cases. To compute $p_{\varepsilon}$, we consider the same $100$ test load instances as in Sec. VI-C. For each load instance, we generate $250$ samples using DiffOPF and count the number of samples that are $\varepsilon$-close to the solution obtained by the conventional interior-point method. The empirical probability $p_{\varepsilon}$ is computed as the fraction of $\varepsilon$-close samples among the total $100\times 250$ generated samples.
Table II reports the theoretical lower bound $M$ on the number of samples required to obtain at least one $\varepsilon$-close solution with confidence level $\delta$, together with the empirical number $N_{\varepsilon}$ of $\varepsilon$-close samples observed among the $M$ generated samples. Comparing $N_{\varepsilon}$ with $M$ provides insight into the conservativeness of the theoretical bound. Note that we present the distance $\varepsilon$ in percentage, as the cost structures of the two systems are different.
The lower bound appears conservative for the small PJM $5$-bus system, where a relatively large fraction of the $M$ samples are $\varepsilon$-close to the optimal solution (i.e., $N_{\varepsilon}$ is often comparable to $M$). In contrast, for the larger IEEE $118$-bus system the bound becomes noticeably tighter, with $N_{\varepsilon}$ remaining small relative to $M$. As $\varepsilon$ increases, fewer samples are required to obtain an $\varepsilon$-close solution. Even for a small tolerance such as $\varepsilon=0.5\%$, DiffOPF requires only a modest number of samples, not exceeding $269$.
@@ VII Conclusion
For future work, we plan to extend the proposed framework beyond warm start applications by explicitly incorporating OPF constraints within the diffusion sampling process. Building on our prior work in [10], where power flow constraints were enforced through gradient guidance, we aim to integrate OPF constraints to ensure the feasibility of generated solutions. In addition, we intend to address the runtime gap between DiffOPF, which takes at most $2.3$ seconds for the IEEE $118$-bus system, and single-valued deep learning OPF solvers that achieve millisecond runtimes. To this end, we will explore accelerating DiffOPF using diffusion distillation techniques, which compress the sampling process into only a few denoising steps [16].